OpenStudy (yamyam70):

2 cos a cos B is equal to what identity?

OpenStudy (anonymous):

\[\cos(a+b) + \cos(a-b)\]

OpenStudy (yamyam70):

I mean like this, 2 cos α cos β = 2 (_________)

OpenStudy (anonymous):

yea.. thats what i meant.. \[2cosAcosB= \cos(A+B) +\cos(A-B)\]

OpenStudy (yamyam70):

what would i put on my blank?

OpenStudy (anonymous):

you think and tell for eg.. \[3 = \frac{1}{2}( ....)\] what will be in the blank!?

OpenStudy (anonymous):

in the blank cos(a+b)+cos(a-b)

OpenStudy (yamyam70):

have no idea. the equsation on my book is indded what you have said, but my online assignment is asking for thw blank

OpenStudy (anonymous):

yea i know.. can you think? :) first answer my example question

OpenStudy (yamyam70):

2 cos a cos b = cos(a+b)+cos(a-b) 2 cos a cos b = 2 (____________ ) = cos(a+b)+cos(a-b)

OpenStudy (anonymous):

yea.. you are so close.. :P.. if u answer my question.. u will be able to answer ur question too!

OpenStudy (yamyam70):

is it 2 ?

OpenStudy (debbieg):

If \(2X=Y\) then \(2X=2\left(\dfrac{Y}{???}\right)\)

OpenStudy (yamyam70):

x ?

OpenStudy (anonymous):

\[5\ apples = \frac{10\ apples}{???}\]

OpenStudy (yamyam70):

2

OpenStudy (anonymous):

so similarly if i said \[3\ apples = \frac{(???)\ apples}{3}\]

OpenStudy (yamyam70):

9 apples?

OpenStudy (anonymous):

put nine.. there.. and tell me is that right answer? :P

OpenStudy (yamyam70):

9/3 = 3

OpenStudy (anonymous):

so it is..correct :) now similarly i can say \[3X = \frac{?X}{2}\]

OpenStudy (yamyam70):

6 :)

OpenStudy (anonymous):

and finally think in similar manner and tell me \[\cos(a+b)+\cos(a-b) = 2[(??) \cos(a+b) + \cos(a-b)]\]

OpenStudy (yamyam70):

1/ 2 !

OpenStudy (anonymous):

Bingo!..

OpenStudy (yamyam70):

thank you!

OpenStudy (anonymous):

your welcome :)!

OpenStudy (yamyam70):

for the patience :) :D

OpenStudy (anonymous):

no problem.. i am a teacher.. i have tons of patience :D :D

OpenStudy (yamyam70):

:)))

OpenStudy (debbieg):

Amen to that, @mashy! :) Just make sure you have one more set of parens in that expression: \(\Large \cos(a+b)+\cos(a-b) = 2[(??) \color{red}(\cos(a+b) + \cos(a-b)\color{red})]\) :)

OpenStudy (yamyam70):

Yup I made sure, thanks debbieG :)